# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ Appendix B Detailed Tables Referenced in Section 9
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
data-defined records. See notes to Table 13 for details of the universe for nonmodal data-defined
@@
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
data-defined records. See notes to Table 13 for details of the universe for modal data-defined
@@ Appendix C Details of the Swapping Experiments in the Main Text
Appendix D The Effects of the 1980 Suppression Rules if Applied to the 2010 Census Summary File 1
To study the suppression rules as implemented in the 1980 Census, we had to coarsen the 63-category race feature. Using the full set of race categories would have resulted in suppressing nearly all tables. We implemented coarsening based on the [OMB:1997, ] version of Statistical Policy Directive 15 as adopted in 1997 by the Department of Justice Voting Section at that time for use in the redistricting that followed the 2000 Census. The coarsened race categories are:
@@ Appendix D The Effects of the 1980 Suppression Rules if Applied to the 2010 Census Summary File 1
  • Table Suppression: Whole tables were suppressed (not published) for geographies with between 1 and 14 persons in any of the race/ethnicity groups; applied to

### v2

@@ Appendix B Detailed Tables Referenced in Section 8
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
data-defined records. See notes to Table LABEL:tab:reid_blockpop_nm for details of the universe for nonmodal data-defined
@@
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
data-defined records. See notes to Table LABEL:tab:reid_blockpop_nm for details of the universe for modal data-defined
@@ Appendix C Details of the Swapping Experiments in the Main Text
Appendix D The Effects of the 1980 Suppression Rules if
Applied to the 2010 Census Summary File 1
To study the suppression rules as implemented in the 1980 Census, we had to coarsen the 63-category race feature. Using the full set of race categories would have resulted in suppressing nearly all tables. We implemented coarsening based on the OMB:1997 version of Statistical Policy Directive 15 as adopted in 1997 by the Department of Justice Voting Section at that time for use in the redistricting that followed the 2000 Census. The coarsened race categories are:
@@ Appendix D The Effects of the 1980 Suppression Rules if
  • Table Suppression: Whole tables were suppressed (not published) for geographies with between 1 and 14 persons in any of the race and ethnicity groups; applied to